OpenStudy (anonymous):

Solve the following system. x − y = 10 x + y = 6 a. (−2, 8) b. (−8, 2) c. (8, −2) d. (8, 2)

OpenStudy (anonymous):

well we can solve it by using different ways! first of all: we have x - y = 10 so x = y + 10

OpenStudy (anonymous):

ok?

OpenStudy (anonymous):

ok

OpenStudy (texaschic101):

just add them, the y's will cancel out and you can solve for x

OpenStudy (anonymous):

then plug in the x value and solve for y

OpenStudy (texaschic101):

exactly :)

OpenStudy (anonymous):

we have x = y +10 so we can put y + 10 + y = 6 :)

OpenStudy (anonymous):

and @texaschic101 told the other way

OpenStudy (texaschic101):

true..you can solve by substitution as well

OpenStudy (anonymous):

ok but I still am confused what is the answer

OpenStudy (anonymous):

try out the subtracting method and see what you get

OpenStudy (anonymous):

ok I got d

OpenStudy (anonymous):